OpenStudy (mouth4war):

What is the expression in radical form? (3p^3q) 3/4

OpenStudy (mathmale):

For added clarity, please type or draw this expression in as\[(3p^3q)^{\frac{ 3 }{ 4 }}\]

OpenStudy (mathmale):

Notice how the (3/4) is a superscript; it's above the level of (3p^3*q). From that (3/4), we see that the "index of the root" is 4 and that the remaining power is a 3. Rule:\[(a^b)^c=a ^{bc}\]

OpenStudy (mathmale):

Another rule relevant here is:\[a ^{\frac{ b }{ c }}=\sqrt[c]{a^b}\]

OpenStudy (mathmale):

The latter rule will govern your conversion of the given expression into radical form.

OpenStudy (mathmale):

Let me try again:\[a ^{\frac{ b }{ c }}=\sqrt[c]{a^b}\]